19 red 11 Blue 20 Yellow given the number of red, blue, and yellow balls in...

19 red

11 Blue

20 Yellow

given the number of red, blue, and yellow balls in a box. In how many ways can one yellow and two red balls be taken from the box, with replacement?

please with the formula
Answer #1

Total balls = 19 + 11 + 20 = 50

P(yellow) = 20/50 = 0.4

P(red) = 19/50 = 0.38

one yellow and two red balls = 20C1 * 19C2

                                                   = (20! / (1! * 19!)) * (19! / (2! * 17!))

                                                   = 20 * 171

= 3420